Quadrilateral ABCD is a parallelogram with AB = 3y – 6, BC = 5x – 8, CD = 2y + 4, and AD = 4x –1. Find x.
Answer by ikleyn(53765)

.
Quadrilateral ABCD is a parallelogram with AB = 3y – 6, BC = 5x – 8, CD = 2y + 4, and AD = 4x –1. Find x.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Parallelogram has opposite sides congruent. i.e. of equal length.

So, you have these equations:

1)  |AB| = |CD|,  or  3y - 6 = 2y + 4.

                  The solution is 3y - 2y = 4 + 6,  or  y = 10.

and


2)  |BC| = |AD|,  or 5x-8 = 4x - 1.

                  The solution is  5x - 4x = 1 + 8,  or  x = 9.

Solved.